# Investigating the Fermi-Hubbard model by the Tensor-Backflow method

### v3

@@ II.2 Optimization method
The wave-function is firstly optimized by the VMC, then further optimized by a Lanczos step. The energy and the $\alpha$-th parameter’s gradient are evaluated through the Markov-Chain-Monte-Carlo (MCMC) process Hubbard_fPEPS ; PEPS_liu :
@@
where the local energy is $E_{\mathrm{loc}}(\mathbf{S})=\sum_{\mathbf{S}^{\prime}}\frac{w(\mathbf{S}^{\prime})}{w(\mathbf{S})}\langle\mathbf{S}^{\prime}|\hat{H}|\mathbf{S}\rangle$, the $O_{\mathrm{loc}}(\mathbf{S})^{\alpha}=\frac{1}{w(\mathbf{S})}\frac{\partial w(\mathbf{S})}{\partial\alpha}$, and $\langle\cdots\rangle$ denotes the average on MCMC samples.
The variational parameters are updated according to the gradient descent method. Here we only adopt the first-order gradient descent due to the low optimization difficulty of the tensor representation. Because of the limited MC sample number, we take the sign of the gradient and apply a constant step size $\delta$: $\alpha^{\prime}=\alpha-\delta\mathrm{sgn}(G^{\alpha})$. Such parameter updating scheme has been successfully used in optimizing high dimensional tensors like the PEPS tJ_fPEPS ; PEPS_liu .
A Lanczos step further improves the representation ability of wave-function $|\Psi_{p=0}\rangle$ by considering an additional wave-function $|\Psi_{p=0}^{\perp}\rangle$ orthogonal to $|\Psi_{p=0}\rangle$ lanczos :
@@
where $A$ and $B$ are parameters to be determined, and $|\Psi_{p=0}\rangle$ is the wave-function obtained after the VMC. The orthogonal wave-function is built by $|\Psi_{p}^{\perp}\rangle=(\hat{H}-E_{p})/\sigma_{p}|\Psi_{p}\rangle$, where the energy expectation $E_{p}=\langle\Psi_{p}|\hat{H}|\Psi_{p}\rangle$ and the variance $\sigma_{p}^{2}=\langle\Psi_{p}|(\hat{H}-E_{p})^{2}|\Psi_{p}\rangle$.
@@ II.3 Computational complexity
For the Tensor-Backflow wave-function, the complexity for generating a coefficient $w(\mathbf{S})$ is $\mathcal{O}(N^{3}Q)$ based on Eq.(3) and Eq.(4), where $\mathcal{O}(N^{3})$ is from the Slater determinant and $\mathcal{O}(Q)$ is from the summation in Eq.(4).
When considering backflow terms from nearest-neighbor (NN) sites, there is QQ=5. The representation ability can be improved by considering backflow terms in the range of next-nearest-neighbor (NNN) sites, there is QQ=9. We denote the above two types as NN and NNN backflows, and the complexities are $\mathcal{O}(N^{3})$ for both backflow types. The maximal representation ability is achieved by considering backflow terms from all sites: $Q=M$, meanwhile the complexity is $\mathcal{O}(MN^{3})$.
When generating a coefficient for the wave-function after a Lanczos optimization, the complexity is $\mathcal{O}(N^{4}Q)$ for generating a coefficient $w(\mathbf{S})$ based on Eq.(6). During the Lanczos optimization, because of the term $\langle\Psi_{p=0}|\hat{H}^{2}|\Psi_{p=0}\rangle$ when calculating the variance $\sigma_{p}^{2}$, the complexity is $\mathcal{O}(N^{5}Q)$.
@@ III Numerical results
where $t$ and $t^{\prime}$ are strengths for NN and NNN hoppings, respectively. $U$ is the strength of on-site interactions, $\hat{c}_{i\sigma}^{\dagger}(\hat{c}_{i\sigma})$ creates (destroys) a particle of spin $\sigma$ on the $i$-th site, and the particle number operator $\hat{n}_{i\sigma}=\hat{c}_{i\sigma}^{\dagger}\hat{c}_{i\sigma}$. We set tt=1 through our investigations. Energy results achieved by wave-functions before and after a Lanczos optimization are evaluated with the MC sample number in the magnitude of $\mathcal{O}(4.5\times 10^{4})$.
@@ III.1 Optimization details
In VMC optimizations, sample number and step size are important for achieving fully convergences. To identify these parameters, we have tested two optimization strategies:
@@
    Under the sample number in the magnitude of $\mathcal{O}(4.5\times 10^{4})$, firstly setting the step size as $2\times 10^{-3}$ and then reducing it to $2\times 10^{-4}$;
@@
    Under the sample number in the magnitude of $\mathcal{O}(4.5\times 10^{4})$, firstly setting the step size as $2\times 10^{-3}$, then increasing the sample number to $\mathcal{O}(4.5\times 10^{5})$ and maintaining the step size, finally reducing the step size to $2\times 10^{-4}$.
Comparisons of two optimization strategies are depicted in Fig.(1). The case in the figure is nn=0.875, UU=8, $t^{\prime}$=0 on the $4\times 16$ lattice under open boundary condition (OBC), with NN backflow terms in the wave-function. From the figure, both strategies converge to similar energies. However, strategy 2 requires less optimization steps. Both strategies achieve similar ground state patterns after fully convergences, and comparisons of ground state patterns are detailed in Appendix.(A). Furthermore, we find that strategy 2 is beneficial to the convergence of all-site backflow, due to the much larger MC sample number. Therefore, strategy 2 is used for optimizing all-site backflow wave-functions. On lattices larger than 64 sites, for both optimization strategies, no more than $2\times 10^{4}$ optimization steps are used for fully convergence.
Figure 2: Energy extrapolations and comparisons respect to fPEPS Hubbard_fPEPS under OBC for the lattice of $8\times 8$(a) and $4\times 16$(b). All cases are for nn=0.875, UU=8 and $t^{\prime}$=0. Energy extrapolation for the lattice of $4\times 16$ under PBC(c) yields to the neural Pfaffian result Hubbard_NN4 . $\sigma^{2}$ is the energy variance. Red and blue solid dots depict Tensor-Backflow with NN and all-site backflow terms. For solid dots of the same color, low and high energies depict results with and without one Lanczos optimization.
@@ III.2 Comparisons with previous results
Table 1: Energy comparisons respect to state-of-the-art results for various filling number $n$ and interaction strength $U$. All cases are for $t^{\prime}$=0. A-PBC denotes the boundary condition of antiperiodic along the x-boundary and periodic along the y-boundary. $E_{p=0}$ and $E_{p=1}$ are energies achieved by Tensor-Backflow without and with one Lanczos optimization, respectively.
In our optimizations, the initial state of Tensor-Backflow is unrestricted Hartree-Fock (UHF) state represented by a tensor of the dimension $[M,N,2]$, where $M$ is the site number $N$ is the particle number and 2 is because of spin. From the initial state, we expand the tensor’s dimension to Eq.(2) to include backflow terms tensorbackflow .
In Fig.(2)(a)(b), red and blue solid dots denote NN and all-site backflow terms, respectively. High and low energy values for same-colored solid dots denote without and with one Lanczos optimization, respectively. Therefore, a Lanczos optimization significantly improves energy precision for either NN backflow and all-site backflow. Comparing obtained energies after one Lanczos optimization and the fPEPS with DD=20, the NN backflow can achieve the energy precision competitive to the SU optimized fPEPS, meanwhile the all-site backflow can achieve the energy precision competitive to GO optimized fPEPS. The Tensor-Backflow has no limitations on boundary conditions, and energies on the $4\times 16$ lattice under PBC are depicted in Fig.(2)(c).
@@
To obtain the energy extrapolation under the limit of zero energy variance, we linearly fit the four solid dots for each lattice size Hubbard_TNS ; lanczos . In Fig.(2), energy extrapolations are denoted by blue solid lines. From Fig.(2)(a)(b), extrapolated energies under OBC for $8\times 8$ is -0.6993, and for $4\times 16$ is -0.6826. Extrapolated energies are close to state-of-the-art energies: -0.6993 for $8\times 8$ and -0.6830 for $4\times 16$, achieved by the fPEPS with DD=20 within GO optimized. Based on results of OBC, the extrapolated energy of PBC in Fig.(2)(c) is -0.7642, matching the state-of-the-art neural Pfaffian result Hubbard_NN4 .
@@
Energy comparisons respect to fPEPS Hubbard_fPEPS , NQS Manybody_benchmark , tensor networks Hubbard_TNS and AFQMC AFQMC_n1 ; AFQMC_sym are detailed in Tab.(1). In the table, $E_{p=0}$ is from the Tensor-Backflow and $E_{p=1}$ is from a Lanczos optimization of the converged wave-function. For OBC, reference energies are from GO optimized fPEPS with DD=20 Hubbard_fPEPS . On lattices of $8\times 8$ and $4\times 16$ under OBC, the wave-function’s type is all-site backflow. On the $16\times 16$ lattice under OBC, the wave-function’s type is NN backflow.
@@
In this work, the relative error of energy $E$ is defined as: $(E-E_{\mathrm{ref}})/|E_{\mathrm{ref}}|$. Under OBC, relative errors of $E_{p=1}$ respect to fPEPS results are $1.6\times 10^{-3}$, $1.3\times 10^{-3}$ and $4.1\times 10^{-4}$ for $8\times 8$, $4\times 16$ and $16\times 16$ lattices, respectively. Therefore, Tensor-Backflow has high precision for OBC. For boundary conditions other than OBC, from the VMC method in literature Manybody_benchmark , there is $E_{\mathrm{ref}}=-0.7544$ for nn=0.875 and UU=8 on the $16\times 16$ lattice under PBC, and $E_{p=1}$ is $1\times 10^{-3}$ lower. For nn=0.875 and UU=10, with antiperiodic along the x-boundary and periodic along the y-boundary, denoted by A-PBC in Tab.(1), the reference energy is from the energy extrapolation result of the tensor network method in literature Hubbard_TNS , meanwhile $E_{p=1}$ is $2.2\times 10^{-3}$ lower. In Tab.(1), for each reference energy from literatures AFQMC_n1 ; AFQMC_sym , $E_{p=1}$ is competitive and the relative error is in the magnitude of $\mathcal{O}(10^{-3})$.
@@
From comparisons in Tab.(1), the Tensor-Backflow can achieve competitive results respect to mainstream methods. Furthremore, we compare our results to recent state-of-the-art NQS method Hubbard_NN3 in Tab.(2). In the table, except the result on OBC (detailed in Sec.III.3), each result was achieved by the first-order-gradient optimization with the step number less than $2\times 10^{4}$, without any pinning-field appied during optimizations. However, for NQS, each optimization was performed by an improved stochastic reconfiguration method with the training procedure consisting of pre-training, applying pinning-fields and the optimization step number in the magnitude of $\mathcal{O}(10^{5})$ Hubbard_NN3 . In particular, for the case of $t^{\prime}$=-0.2, nn=0.875, UU=8 on the $12\times 12$ lattice under PBC, the achieved energy is $8.1\times 10^{-4}$ lower than the NQS result, and the optimization is performed by first-order-gradient method in less than $1\times 10^{4}$ steps, optimization procedure is detailed in Sec.III.3. Therefore, the Tensor-Backflow is efficient for optimizations meanwhile has a strong state representation ability.
Tab.(3) denotes parameter numbers used in Tensor-Backflow for various lattice sizes and backflow types, as well as energy precision respect to two state-of-the-art methods under OBC. According to Eq.(2), for the filling of nn=0.875 on 64-site lattices, the parameter number for NN backflow and all-site backflow are 286720 and 3670016, respectively. Meanwhile, the parameter number for NN backflow on the $16\times 16$ lattice is 4587520. In the table, we compare $E_{p=1}$ to results achieved by state-of-the-art fPEPS Hubbard_fPEPS and NQS Hubbard_NN3 . From energy comparisons in the table, NN backflows can achieve low energy relative errors in the magnitude of $\mathcal{O}(10^{-3})$. Meanwhile, considering all backflow terms can further improve energy precision.
The parameter number of Tensor-Backflow is much lower comparing to the state-of-the-art fPEPS. On the $16\times 16$ lattice under OBC, the fPEPS with DD=14 has $5.7\times 10^{6}$ parameters, and the achieved energy is -0.7224. The fPEPS with DD=16 has $9.4\times 10^{6}$ parameters, and the achieved energy is -0.7251. The fPEPS with DD=18 has $1.5\times 10^{7}$ parameters, and the achieved energy is -0.7260. The fPEPS with the maximal bond dimension DD=20 has $2.2\times 10^{7}$ parameters, with the similar energy as DD=18 Hubbard_fPEPS . For the Tensor-Backflow on the $16\times 16$ lattice, there are $E_{p=0}=$-0.7219 and $E_{p=1}=$-0.7257. Under the similar energy precision, without a Lanczos optimization, the Tensor-Backflow has many lower parameters comparing to the fPEPS with DD=14. After a Lanczos optimization, same energy precision is achieved by the Tensor-Backflow with significantly lower parameters comparing to the fPEPS with DD=18,20. Therefore, the Tensor-Backflow is an efficient state representation.
@@
The filling of nn=0.875 is important for understanding the mechanism of superconductivity, especially when the value of $t^{\prime}$ is non-zero Hubbard_SC1 . The ground energies are highly competitive between various methods Hubbard_stripe . In this subsection, we present more details such as ground state properties and structure factors when nn=0.875 for both $t^{\prime}$=0 and $t^{\prime}$=-0.2. We firstly investigate results achieved by Tensor-Backflow when $t^{\prime}$=0.
In this work, most results from Tensor-Backflow are achieved by direct optimizations. For a direct optimization, an UHF is optimized beforehand, followed by constructing the Tensor-Backflow through expanding the tensor’s dimension to add backflow terms, meanwhile the Tensor-Backflow is optimized using either strategy from Sec.III.1. There are no pinning-fields applied during optimizations. When optimizing the UHF, the interaction strength is denoted as $U_{\mathrm{UHF}}$. Unless specifically emphasized, there is $U_{\mathrm{UHF}}=U$, and the UHF is not sufficiently optimized when $U_{\mathrm{UHF}}=U$ to avoid introducing any prior bias UHF . Futhermore, Tensor-Backflow can override properties from the initial UHF, namely, the converged wave-function is essentially a different state from the initial UHF, detailed comparisons are depicted in Appendix.(B).
Fig.(3)(a)(b) depict the linear stripe order of the ground state for UU=8 on the $16\times 16$ lattice under PBC by a direct optimization. Fig.(3)(a) depicts the average spin per site: $s_{i}^{z}=(n_{i}^{\uparrow}-n_{i}^{\downarrow})/2$ and Fig.(3)(b) depicts the average density per site: $n_{i}=n_{i}^{\uparrow}+n_{i}^{\downarrow}$. As depicted by the ground state’s pattern, periods for the SDW and CDW are 16 and 8, respectively. For PBC, the ground state should be translational invariant thus an obvious pattern is not consistent with the true ground state. Since translational invariance is not enforced in the variational wave-function, even the obtained energy is competitive, the optimization is converged to a subplace of the true ground state.
@@
Table 4: Energy comparisons of $E_{p=1}$ between different initializations for nn=0.875 and $t^{\prime}$=0, with NN backflow in wave-functions. For the direct optimization of each case, the initial UHF is insufficiently optimized with $U_{\mathrm{UHF}}=U$. The direct optimization of UU=8 achieves the ground state of linear stripe order, and such order is preserved when it is transferred for other $U$ and boundary conditions. Energy convergences of transferring from UU=8 are detailed in Appendix.C. Energies with linear stripe orders are much lower than those by direct optimizations.
@@
For nn=0.875, UU=8, $t^{\prime}$=0 on the $16\times 16$ lattice under OBC, we transfer the wave-function converged under PBC as the initial state and then performed optimizations. The achieved state-of-the-art energy is -0.7257 as reported in Tab.(1) and (3). Meanwhile, Fig.(3)(c)(d) depict the average spin and charge density per site, respectively. Even the stripe order in the initial state has the same period as the ground state under OBC, the optimization can slightly modify the stripe’s pattern according to the boundary condition. Patterns of the ground state under OBC yield to fPEPS results Hubbard_fPEPS .
For other interaction strengths such as UU=10 and 12, Tab.(4) denote energy comparisons between the linear stripe order and direct optimizations, where linear stripe order is achieved when transferring the wave-function converged for UU=8 and PBC. In particular, direct optimization achieves diagonal stripe order when UU=12 under PBC. Ground state patterns of direct optimizations are depicted in Appendix.B. Energy convergences of transferring from UU=8 is detailed in Appendix.C. Energy comparisons are depicted in Tab.(4), and all energies are evaluated by pp=1 wave-functions. Because the energy of linear stripe order is much lower than other orders achieved by direct optimizations, the ground state is linear stripe order for large $U$.

### v4

@@ II.2 Optimization method
We use the VMC method to optimize the wave-function first, followed by a Lanczos step. In the VMC, the energy and the gradient of the $\alpha$-th parameter are estimated through the Markov-Chain-Monte-Carlo (MCMC) process Hubbard_fPEPS ; PEPS_liu :
@@
where the local energy is $E_{\mathrm{loc}}(\mathbf{S})=\sum_{\mathbf{S}^{\prime}}\frac{w(\mathbf{S}^{\prime})}{w(\mathbf{S})}\langle\mathbf{S}^{\prime}|\hat{H}|\mathbf{S}\rangle$, the $O_{\mathrm{loc}}(\mathbf{S})^{\alpha}=\frac{1}{w(\mathbf{S})}\frac{\partial w(\mathbf{S})}{\partial\alpha}$, and $\langle\cdots\rangle$ denotes the average on MC samples.
In the VMC optimization, the variational parameters are updated using the gradient descent method. Since the absolute values of the gradients are difficult to be estimated with finite MCMC samples, we update the parameters based on the sign of the gradient and apply a constant step size $\delta$: $\alpha^{\prime}=\alpha-\delta\mathrm{sgn}(G^{\alpha})$. This parameter update scheme has been successfully applied in optimizing high-dimensional tensors, such as the PEPS tJ_fPEPS ; PEPS_liu .
To extend the representation ability beyond backflow corrections, the wave-function is further optimized using a Lanczos step. In the Lanczos procedure, the wave-function is optimized by constructing Krylov subspaces:
@@
where $A$ and $B$ are parameters to be determined, and $|\Psi_{p=0}\rangle$ is the wave-function optimized by the VMC. The orthogonal wave-function is built by $|\Psi_{p}^{\perp}\rangle=(\hat{H}-E_{p})/\sigma_{p}|\Psi_{p}\rangle$, where the energy expectation and the variance are: $E_{p}=\langle\Psi_{p}|\hat{H}|\Psi_{p}\rangle,\ \sigma_{p}^{2}=\langle\Psi_{p}|(\hat{H}-E_{p})^{2}|\Psi_{p}\rangle$.
@@ II.3 Computational complexity
For the Tensor-Backflow wave-function, the complexity of generating a coefficient $w(\mathbf{S})$ is $\mathcal{O}(N^{3}Q)$, as derived from Eq.(5) and Eq.(6), where $\mathcal{O}(N^{3})$ comes from the Slater determinant and $\mathcal{O}(Q)$ from the summation in Eq.(6).
When considering backflow terms from nearest-neighbor (NN) sites, QQ=5. The representation ability is improved by including backflow terms from next-nearest-neighbor (NNN) sites, where QQ=9. We refer to these two cases as NN and NNN backflows, and the computational complexity is $\mathcal{O}(N^{3})$ for both types. The maximal representation ability is achieved by including backflow terms from all sites, where QQ=$M$, leading to a complexity of $\mathcal{O}(MN^{3})$.
For generating a coefficient after Lanczos step, the complexity is $\mathcal{O}(N^{4}Q)$, based on Eq.(8). During the Lanczos step, the term $\langle\Psi_{p=0}|\hat{H}^{2}|\Psi_{p=0}\rangle$ contributes to the calculation of the variance $\sigma_{p}^{2}$, resulting in a complexity of $\mathcal{O}(N^{5}Q)$.
@@ III Numerical results
where $t$ and $t^{\prime}$ are strengths for NN and NNN hoppings, respectively. $U$ is the strength of on-site interactions, $\hat{c}_{i\sigma}^{\dagger}(\hat{c}_{i\sigma})$ creates (destroys) a particle of spin $\sigma$ on the $i$-th site, and the particle number operator $\hat{n}_{i\sigma}=\hat{c}_{i\sigma}^{\dagger}\hat{c}_{i\sigma}$. We set tt=1 through our investigations. Energy results achieved by wave-functions before and after a Lanczos step are evaluated with the MC sample number in the magnitude of $\mathcal{O}(4.5\times 10^{4})$.
@@ III.1 Optimization details
In VMC optimizations, the number of samples and the step size are crucial for achieving full convergence. To determine the optimal values for these parameters, we tested two optimization strategies:
@@
    With a sample size of $\mathcal{O}(4.5\times 10^{4})$, the step size is initially set to $2\times 10^{-3}$ and then reduced to $2\times 10^{-4}$.
@@
    With a sample size of $\mathcal{O}(4.5\times 10^{4})$, the step size is initially set to $2\times 10^{-3}$, the sample size is then increased to $\mathcal{O}(4.5\times 10^{5})$ while maintaining the step size, and finally, the step size is reduced to $2\times 10^{-4}$.
The comparison between these two strategies is shown in Fig.(1). The specific case depicted in the figure is nn=0.875, UU=8, $t^{\prime}$=0 on a $4\times 16$ lattice under open boundary conditions (OBC), with NN backflow terms in the wave-function. From the figure, both strategies converge to similar energies. However, strategy (2) requires fewer optimization steps. Both strategies yield similar ground state patterns after full convergence, and further comparisons of the ground state patterns are provided in App. (A). Additionally, we find that strategy (2) is more effective for the convergence of all-site backflow due to the significantly larger number of MCMC samples. Therefore, strategy (2) is used for optimizing wave-functions with all-site backflow. For lattices larger than 64 sites, both strategies require no more than $2\times 10^{4}$ optimization steps to achieve full convergence.
Figure 2: Energy extrapolations and comparisons respect to fPEPS Hubbard_fPEPS under OBC for the lattice of $8\times 8$(a) and $4\times 16$(b). All cases are for nn=0.875, UU=8 and $t^{\prime}$=0. Energy extrapolation for the lattice of $4\times 16$ under PBC(c) yields to the neural Pfaffian result Hubbard_NN4 . $\sigma^{2}$ is the energy variance. Red and blue solid dots depict Tensor-Backflow with NN and all-site backflow terms. For solid dots of the same color, low and high energies depict results with and without one Lanczos step.
@@ III.2 Comparisons with previous results
In Fig.(2)(a)(b), red and blue solid dots represent the results for NN and all-site backflow terms, respectively. The higher and lower energy values for each color correspond to results without and with one Lanczos step, respectively. This demonstrates that a single Lanczos step significantly improves the energy precision for both NN and all-site backflow terms. When comparing the energies obtained after one Lanczos step to those from the fPEPS with bond dimension DD=20, we find that the NN backflow achieves energy precision comparable to the SU-optimized fPEPS, while the all-site backflow achieves precision competitive with the GO-optimized fPEPS. The Tensor-Backflow method imposes no limitations on boundary conditions, and energies for the $4\times 16$ lattice under PBC are shown in Fig.(2)(c).
@@
Table 1: Energy comparisons respect to state-of-the-art results for various filling number $n$ and interaction strength $U$. All cases are for $t^{\prime}$=0. A-PBC denotes the boundary condition of antiperiodic along the x-boundary and periodic along the y-boundary. $E_{p=0}$ and $E_{p=1}$ are energies achieved by Tensor-Backflow without and with one Lanczos step, respectively.
@@
To perform energy extrapolation under the limit of zero energy variance, we linearly fit the four solid dots for each lattice size Hubbard_TNS ; lanczos . The extrapolated energies are indicated by blue solid lines in Fig.(2). From Fig.(2)(a)(b), the extrapolated energies under OBC are -0.6993 for the $8\times 8$ lattice and 0.6826-0.6826 for the $4\times 16$ lattice. These extrapolated values are in close agreement with state-of-the-art energies: -0.6993 for the $8\times 8$ lattice and 0.6830-0.6830 for the $4\times 16$ lattice, as achieved by the fPEPS with DD=20 and GO optimization. The extrapolated energy for PBC shown in Fig.(2)(c) is 0.7642-0.7642, which matches the state-of-the-art neural Pfaffian result Hubbard_NN4 .
@@
Energy comparisons with fPEPS Hubbard_fPEPS , NQS Manybody_benchmark , tensor networks Hubbard_TNS , and AFQMC AFQMC_n1 ; AFQMC_sym are provided in Tab.(1). In the table, $E_{p=0}$ represents the energy from the Tensor-Backflow wave-function, and $E_{p=1}$ represents the energy obtained after a Lanczos step on the converged wave-function. For OBC, the reference energies are taken from GO-optimized fPEPS with DD=20 Hubbard_fPEPS . For the $8\times 8$ and $4\times 16$ lattices under OBC, the wave-function corresponds to all-site backflow, while for the $16\times 16$ lattice under OBC, the wave-function uses NN backflow.
@@
In this work, the relative error of energy $E$ is defined as: $(E-E_{\mathrm{ref}})/|E_{\mathrm{ref}}|$. where $E_{\mathrm{ref}}$ is the reference energy. For OBC, the relative errors of $E_{p=1}$ with respect to the fPEPS results are $1.6\times 10^{-3}$, $1.3\times 10^{-3}$, and $4.1\times 10^{-4}$ for the $8\times 8$, $4\times 16$, and $16\times 16$ lattices, respectively. This demonstrates that Tensor-Backflow achieves high precision for OBC. For other boundary conditions, the reference energy from the VMC method in the literature Manybody_benchmark is $E_{\mathrm{ref}}$=-0.7544 for nn=0.875 and UU=8 on the $16\times 16$ lattice under PBC, with $E_{p=1}$ being $1\times 10^{-3}$ lower. For nn=0.875 and UU=10, under antiperiodic along the x-boundary and periodic along the y-boundary, denoted by A-PBC in Tab.(1), the reference energy is obtained from energy extrapolation results of the tensor network method in the literature Hubbard_TNS , with $E_{p=1}$ being $2.2\times 10^{-3}$ lower. In Tab.(1), for each reference energy from the literature AFQMC_n1 ; AFQMC_sym , the energy after one Lanczos step is competitive, with relative errors on the order of $\mathcal{O}(10^{-3})$.
@@
From the comparisons presented in Tab.(1), the Tensor-Backflow method achieves results that are competitive with those from mainstream approaches. Furthermore, we compare our results to the recent state-of-the-art NQS method Hubbard_NN3 in Tab.(2). In the table, all results (except those for OBC, detailed in Sec. (III.3)) were obtained using first-order gradient optimization, with fewer than $2\times 10^{4}$ optimization steps and without applying any pinning fields during the optimization process. In contrast, for the NQS method, each optimization involved an improved stochastic reconfiguration procedure, pre-training, the application of pinning fields, and a large optimization step number Hubbard_NN3 . In particular, for the case with $t^{\prime}$=-0.2, nn=0.875, UU=8 on the $12\times 12$ lattice under PBC, the achieved energy from the Tensor-Backflow method is $8.1\times 10^{-4}$ lower than the corresponding NQS result. Moreover, this optimization was performed using the first-order gradient method in fewer than $1\times 10^{4}$ steps, with the optimization procedure detailed in Sec. (III.3). Therefore, the Tensor-Backflow method not only demonstrates strong representational capabilities but is also efficient in terms of optimization.
Tab.(3) shows the number of parameters used in the Tensor-Backflow method for various lattice sizes and backflow types, along with the corresponding energy precision compared to two state-of-the-art methods under OBC. According to Eq.(4), for a filling of nn=0.875 on 64-site lattices, the parameter numbers for NN and all-site backflow are 286,720 and 3,670,016, respectively. For the $16\times 16$ lattice with NN backflow, the parameter number is 4,587,520. In the table, we compare the energy obtained after one Lanczos step ($E_{p=1}$) with results from the state-of-the-art fPEPS Hubbard_fPEPS and NQS Hubbard_NN3 methods. From the energy comparisons, we observe that NN backflow achieves relative energy errors on the order of $\mathcal{O}(10^{-3})$. Moreover, incorporating all backflow terms further improves the energy precision.
The number of parameters in Tensor-Backflow is significantly lower compared to the state-of-the-art fPEPS method Hubbard_fPEPS . For the $16\times 16$ lattice under OBC, the fPEPS with bond dimension DD=14 contains $5.7\times 10^{6}$ parameters, achieving an energy of 0.7224-0.7224. The fPEPS with DD=16 has $9.4\times 10^{6}$ parameters and achieves an energy of 0.7251-0.7251, while the fPEPS with DD=18 uses $1.5\times 10^{7}$ parameters and yields an energy of 0.7260-0.7260. The fPEPS with the maximum bond dimension, DD=20, requires $2.2\times 10^{7}$ parameters and produces an energy similar to that obtained with DD=18 Hubbard_fPEPS . For Tensor-Backflow on the $16\times 16$ lattice, the energy without a Lanczos step ($E_{p=0}$) is 0.7219-0.7219, and after the Lanczos step ($E_{p=1}$), it improves to 0.7257-0.7257. While achieving comparable energy precision, Tensor-Backflow requires fewer parameters than fPEPS with DD=14 when no Lanczos step is applied. After the Lanczos step, Tensor-Backflow achieves the same energy precision with significantly fewer parameters than fPEPS with DD=18 or DD=20. This demonstrates that Tensor-Backflow is a highly efficient state representation method.
@@
Table 4: Energy comparisons of $E_{p=1}$ between different initializations for nn=0.875 and $t^{\prime}$=0, with NN backflow in wave-functions. For the direct optimization of each case, the initial UHF is insufficiently optimized with $U_{\mathrm{UHF}}=U$. The direct optimization of UU=8 achieves the ground state of linear stripe order, and such order is preserved when it is transferred for other $U$ and boundary conditions. Energy convergences of transferring from UU=8 are detailed in App.(C). Energies with linear stripe orders are much lower than those by direct optimizations.
The filling of nn=0.875 is important for understanding the mechanisms of superconductivity, particularly when the NNN hopping term $t^{\prime}$ is non-zero Hubbard_SC1 . The ground-state energies at this filling exhibit strong competition among various methods Hubbard_stripe . In this subsection, we provide a detailed analysis of the ground-state properties and structure factors for nn=0.875, considering both $t^{\prime}$=0 and $t^{\prime}$=-0.2. We begin by presenting the results obtained from the Tensor-Backflow method for the case of $t^{\prime}$=0.
In this work, most of the results obtained from Tensor-Backflow are achieved through direct optimization. For a direct optimization, a UHF state is initially optimized, followed by the construction of the Tensor-Backflow by expanding the tensor’s dimensions to include backflow terms. The Tensor-Backflow wave-function is then optimized using one of the strategies outlined in Sec. (III.1), with no pinning fields applied during the optimization process. When optimizing the UHF state, the interaction strength is denoted as $U_{\mathrm{UHF}}$. Unless otherwise specified, we set $U_{\mathrm{UHF}}=U$. It is important to note that the UHF is not fully optimized when $U_{\mathrm{UHF}}=U$, in order to avoid introducing prior bias UHF . Furthermore, Tensor-Backflow can override the properties of the initial UHF state, meaning that the converged wave-function is, in essence, a different state from the original UHF. A detailed comparison of the UHF and Tensor-Backflow states is provided in App.( B).
Fig.(3)(a) and (b) show the linear stripe order of the ground state for UU=8 on a $16\times 16$ lattice under PBC, obtained through direct optimization. In Fig.(3)(a), we present the average spin per site, defined as $s_{i}^{z}=(n_{i}^{\uparrow}-n_{i}^{\downarrow})/2$, while Fig.(3)(b) shows the average density per site, $n_{i}=n_{i}^{\uparrow}+n_{i}^{\downarrow}$. From the ground-state patterns, we observe that the periods for the spin-density wave (SDW) and charge-density wave (CDW) are 16 and 8, respectively. For PBC, the ground state should be translational invariant thus an obvious pattern is not consistent with the true ground state. Since translational symmetry is not enforced in the variational wave-function, the optimization may converge to a subspace of the true ground state, even though the obtained energy is competitive.
For nn=0.875, UU=8, $t^{\prime}$=0 on a $16\times 16$ lattice with OBC, we use the wave-function converged under PBC as the initial state and perform further optimizations. The achieved state-of-the-art energy is -0.7257, as reported in Tab.(1) and (3). Meanwhile, Fig.(3)(c) and (d) show the average spin and charge density per site, respectively. Although the stripe order in the initial state has the same period as the ground state under OBC, the optimization process slightly modifies the stripe pattern to accommodate the boundary conditions. The resulting ground-state patterns under OBC are consistent with those from the fPEPS method Hubbard_fPEPS .
For other interaction strengths such as UU=10 and UU=12, Tab.(4) presents energy comparisons between the linear stripe order and the results from direct optimizations. The linear stripe order is obtained by transferring the wave-function converged for UU=8 and PBC. Notably, for UU=12 under PBC, the direct optimization leads to a diagonal stripe order. Ground-state patterns from direct optimizations are shown in App.( B), and the energy convergence of transferring from UU=8 is detailed in App.( C). Energy comparisons are summarized in Tab.(4), with all energies evaluated using the pp=1 wave-function. Since the energy of the linear stripe order is significantly lower than that of other orders achieved by direct optimizations, we conclude that the ground state for large $U$ is the linear stripe order.